OpenStudy (anonymous):

Sandra can paint a kitchen in 3 hours and Charles can paint the same kitchen in 5 hours. How long would it take for both working together to paint the kitchen? a. 4 hour b. 8/15 hr c. 1 7/8 hr d. 8 hr

OpenStudy (mertsj):

Sandy can get 1/3 of the job done each hour. Chuck can get 1/5 of the job done each hour. Let's say that together it takes them x hours. That would mean that together they can get 1/x of the job done each hour. So we have: The part Sandy does + the part Chuck does = part they do together. \[\frac{1}{3}+\frac{1}{5}=\frac{1}{x}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (mertsj):

yw